OpenStudy (anonymous):

There is a cone inside a cube. The diameter of the cone is 4 cm. The length,height,and width of the cube are 6 cm each. What is the volume of the cube not occupied by the cone? ( Use three for pi )

OpenStudy (anonymous):

@amistre64

OpenStudy (raden):

Vcube - Vcone

OpenStudy (anonymous):

thanks. u use volume of rect prism for the cube?

OpenStudy (raden):

Vcone = 1/3 * pi*r^2*t Vcube = side*side*side

OpenStudy (anonymous):

so 36*6-1/3*3*4^2

OpenStudy (anonymous):

wait, the height of the cone is not given

OpenStudy (anonymous):

only the diameter

OpenStudy (raden):

yeah, u didnt mention for heigh of cone ?

OpenStudy (anonymous):

the only thing given is the diameter of the base. the cone is upside down inside the cube

OpenStudy (anonymous):

and the only ,easurement given is the diameter..a line is across the circle/base and says 4 cm

OpenStudy (anonymous):

What do you suppose the height is ?

OpenStudy (raden):

assume that the height of cone same with cube, maybe...

OpenStudy (anonymous):

magbe. and so, what would the equation be?

OpenStudy (raden):

Vcone = 1/3 * pi*(2)^2*4 = 16pi/3 Vcube = 4*4*4 = 64 subtracted from Vcube, gives 64-16pi/3

OpenStudy (anonymous):

thanks. ill calculate now.

OpenStudy (raden):

that's enough as the answer

OpenStudy (anonymous):

is it 48 cm^3?

OpenStudy (anonymous):

ok. thank u

OpenStudy (anonymous):

but if asked for a specific one,would 48 suffice?

OpenStudy (raden):

just convert pi=3.14

OpenStudy (anonymous):

they tell u to use 3 for pi

OpenStudy (raden):

it will easier, so funy... = 64 - 16*3/3 = 64 - 16 = 48 u r right :)